How long does it take to travel 4/10 of a mile at 45mph
Answer by mukhopadhyay(490)

45 miles in 1 hour
=> 1 mile in 1/45 hour
=> 4/10 miles in (4/10)(1/45) = 4/450 = 2/225 hour
= 120/225 minute = .54 minutes (approx)